# Roadside Monocular 3D Detection Prompted by 2D Detection

### v3

@@ Abstract
Figure 1: A summary of our work. (a) To approach roadside monocular 3D detection, we introduce a novel detector design Promptable 3D detector (Pro3D) that exploits 2D detections to prompt the 3D detector for 3D prediction. Pro3D can exploit any 2D detector and 3D detector with minimal modificaitons. (b) The Pro3D design is motivated by the observation that a simple 2D detector DINO [62] significantly outperforms state-of-the-art roadside monocular 3D detectors (e.g., BEVHeight [56] and BEVSpread [48]) w.r.t 2D metrics (on the DAIR-V2X-I dataset). This implies that 2D detection is an “easier” task than monocular 3D detection – with a trained 2D detector, training 3D detector can be “simplified” by learning to lift 2D detections to the 3D space. Moreover, as roadside camera pose is fixed, we derive a backbround image as scene prior, incorporating which remarkably boosts 3D detection performance. (c) As a summary of results, our Pro3D significantly outperforms prior works on the DAIR-V2X-I benchmark. Refer to Table 1 and 2 for comprehensive results.
@@ 1 Introduction
Roadside monocular 3D detection is a practical yet challenging task that requires, over a single RGB frame, detecting objects of predefined classes (e.g., vehicle and pedestrian) and predicting their 3D properties such as the locations, shape, and orientations in bird’s-eye-view (BEV) [56]. It has broad applications [61, 59] in vehicle-vehicle communication [9, 17], vehicle-infrastructure cooperative perception [1], and intelligent traffic control [33, 47].
@@
Status quo. Roadside monocular 3D detection is an ill-posed problem as it requires inferring 3D information from 2D cues. Fortunately, camera pose is fixed on roadside infrastructure, making it possible to learn from data to infer 3D properties from a 2D image. To foster the research of roadside monocular 3D detection, the community has contributed large-scale benchmark datasets, such as DAIR-V2X-I [61] and Rope3D [59]. Existing methods train sophisticated deep neural networks in an end-to-end fashion [19, 18, 56, 48]. For monocular 3D detection, a notorious difficulty is precisely estimating depth for objects [52, 38, 12]. To mitigate this difficulty, prior works propose to convert depth estimation to object height estimation and exploits camera parameters to infer object depth [56], or consider scene geometry to carefully pool features for better 3D prediction [48].
@@
Motivation. We observe that the state-of-the-art roadside 3D detectors BEVHeight [56, 58] and BEVSpead [48] significantly underperform 2D detectors (e.g., DINO [62]) w.r.t 2D detection metrics, as shown by the precision-recall curves on the DAIR-V2X-I dataset [61] in Fig. 1b. Here, we train the 2D detector DINO [62] on this dataset, using the generated 2D bounding boxes by projecting 3D cuboid annotations onto the image plane. To evaluate BEVHeight w.r.t 2D detection metrics, we project its 3D detections onto the image plane as its 2D detections. This comparison has several important implications: (1) 2D detection is an “easier” task than monocular 3D detection, and (2) the fixed camera pose or scene geometry provides an informative prior that greatly simplifies 2D detection. The results motivate us to exploit 2D detections and scene priors to facilitate 3D detector training, leading to our novel detector design (Fig. 1a), Promptable 3D Detector (Pro3D).
@@
Technical insights. Training a 3D detector involves optimizing multiple objectives, such as 2D box coordinate regression, depth estimation, orientation regression, and object classification. Intuitively, a compound holistic objective that optimizes them all together might cause the missing detection of some hard objects as an attempt to handle the difficulty in regressing coordinates, depth and orientation. In contrast, training a 2D detector is simpler that incorporates fewer loss terms, i.e., only a 2D box coordinate regression loss and a classification loss. 2D detection is also well-studied in the literature that develops a multitude of useful techniques. These help a 2D detector outperform 3D detectors on 2D metrics, even when both are trained on the same data. This observation leads to our core insight that exploits a 2D detector to facilitate 3D detector training (Fig. 1a). Intuitively, 2D detections pinpoint the target objects in the image and the 3D detector can be simply thought of as learning to lift them to 3D BEV. From a contemporary view, our approach uses a 2D detector to prompt 3D detector, hence we name our framework Promptable 3D Detector (Pro3D). Moreover, as the camera is fixed in roadside perception, we derive a “blank” scene background as a scene prior from training frames (Fig. 4). We incorporate the scene prior as a part of input to the 3D detector. Our experiments validate the usefulness of exploiting the scene prior, and importantly, the design of Pro3D for roadside monocular 3D detection.
@@
    To better solve roadside monocular 3D detection, we present an embarrassingly simple method to obtain a scene prior: detecting and removing objects in frames, and averaging the pixels across frames towards a “blank” scene background as the scene prior. Incorporating the scene prior greatly enhances 3D detection.
@@ 2 Related Work
Roadside monocular 3D detection helps smart transportation and smart city by enlarging the range of perception with a camera hung on a high infrastructure [56]. To foster research of roadside 3D detection, the community established large-scale datasets such as DAIR-V2X-I [61] and Rope3D [59]. As the roadside camera is typically fixed, multiple methods propose to exploit this prior knowledge to approach in roadside 3D detection [14, 40, 56, 58, 48]. For example, BEVHeight [56] trains a model to regress towards object height instead of object distance or depth, as done previously [18] which finds that regressing height is an easier task than regressing depth given the fixed camera pose. BEVSpread [48] further considers the position approximation error in the voxel pooling process based on BEVHeight. Different from existing methods, we identify a fruitful space to improve existing methods (Table 1) by exploiting a 2D detector (Fig. 1b) to produce 2D detections, using which to facilitate 3D detection (Fig. 1a). Further, we derive a scene prior by exploiting the fixed camera pose, and incorporate the scene prior to enhance 3D detection.
@@
2D object detection aims to detect objects of predefined categories in a 2D image [10, 20]. Prevailing approaches train deep neural networks, which typically consist of a backbone to encode input images, and a detector-head for box regression and classification [36, 35, 22, 44]. Recently, using the transformer architecture significantly improves 2D detection [8, 65, 62]. In this work, we explore how to leverage a 2D detector [62, 44] to train better 3D detectors on roadside scenes. It is important to note that any 2D detectors can be used in our framework Pro3D.
@@
Figure 2: We study three methods of prompt encoding and fusion, used to exploit a 2D detector to facilitate 3D detection. Design (a) concatenates feature maps extracted by the 2D detector and the 3D detector’s backbone. Design (b) extracts a feature vector based on a 2D detection’s coordinates, encodes it through a prompt encoder (implemented as an MLP), and attentively fuses it with the feature map extracted by the 3D detector’s backbone. Design (c) encodes a 2D detection, a 5-dim vector (coordinates $x$ and $y$, object width $w$ and height $h$, and the predicted class label) as the prompt, and attentively fuses the encoded detection with the feature map of the 3D detector’s backbone through the fusion module. Somewhat surprisingly, the third performs the best (Table 6)!
@@
Prompting, enabled by large language models (LLMs), is first recognized in Natural Language Processing (NLP) [6, 39, 51]. Briefly, given an LLM, using a text prompt can instruct the LLM to produce desired text outputs. Inspired by this, the computer vision community endeavors to train large visual/multimodal models that can take text prompts to generate, restore and edit images [32, 23, 29]. Some works use images as prompts to accomplish vision tasks such as segmentation and inpainting [2, 49]. Beyond visual and text prompts, some recent works use more diverse information as prompts to guide models to produce desired output. For example, the Segment Anything Model (SAM) accepts bounding boxes, segmentation masks, and points as prompts towards interactive segmentation [15]. In this work, we propose the promptable 3D detector design (Pro3D), which trains a 3D detector that accepts prompts given by a pre-trained 2D detector. The prompt can be as simple as 2D detections, i.e., predicted box coorediates and class labels. Our Pro3D greatly boosts roadside monocular 3D detection performance (Fig. 1c, Table 1).
@@ 3 Promptable 3D Detector
Problem definition. Roadside monocular 3D detection requires detecting objects in a 2D RGB image and predicting their 3D information $(x,y,z,w,h,l,yaw)$, where $(x,y,z)$ is the object’s center location in the 3D world, $(w,h,l)$ denotes the width, height and length of the cuboid capturing the 3D object, and $yaw$ denotes the orientation of the cuboid on the BEV plane.
@@ 3.1 The Proposed Framework: Pro3D
Our Pro3D exploits a 2D detector for 3D detection and has two core modules to connect the two detectors (Fig. 1a): a prompt encoder and a fusion module. The former encodes 2D detections as new features to prompt the 3D detector. The latter attentively fuses the encoded prompts with the 3D detector backbone for 3D detection. The 3D detector head makes 3D predictions based on the fused features. We elaborate on the two modules below, followed by scene prior generation.
@@
Remark 1: Using 2D detections as prompts yields better 3D detection performance than using features. We explain this by first noting different features extracted from a 2D detector: (1) global feature maps, and (2) local features at a specific position of the feature maps. Global feature maps contain all information from the input image, beneficial to object localization but lacking salient signals for semantic prediction. In contrast, local features may be informative for semantic prediction but may not contain spatial information for localization. Different from these features, the simplistic 2D detections (consisting of box coordinates, object sizes, and predicted labels) can well localize target objects in the input image and deliver their well-predicted labels. By explicitly incorporating the 2D detections, a 3D detector can be trained to lift them to the 3D space towards 3D detections. This simplifies the task of roadside monocular 3D detection and facilitates 3D detector training.
@@
Remark 2: Different from existing methods, our Pro3D explicitly leverages 2D detections. Leveraging 2D box annotations to facilitate 3D detector training is a de facto practice. One recent work similar to ours is MV2D [50], which trains a 3D detector from scratch for joint 2D and 3D detections. It is important to note that MV2D fuses RoI features extracted from the 2D detector with those of the 3D detector’s backbone for 3D detection. This is similar in spirit to our explored baseline method (Fig. 2b). Differently, our final method Pro3D explicitly leverages a 2D detector, without extracting its features, to facilitate 3D detection. Our experiments demonstrate that our Pro3D significantly outperforms MV2D (Table 1). Importantly, owing to the modular design, Pro3D can readily exploit any 2D detectors – using more advanced ones yields better 3D detection performance (Fig. 6).
Fusion module fuses the encoded prompt and the feature maps of the 3D detector backbone (Fig. 2). If the prompt is represented by a feature map extracted from the 2D detector, a straightforward fusion method is to concatenate these feature maps (Fig. 2a), as widely done in the literature [25, 31]. Yet, our final fusion module fuses feature vectors which represent 2D detections as prompts. To fuse the feature vectors, we take inspiration from the Transformer architecture [43, 15] and adopt an attention module. Refer to Fig. 3 for an illustration diagram. Concretely, for an input image $X$ which goes through the 3D detector backbone $F$, its feature map is denoted as $F(X)$. The prompt encoder $F_{prompt}$ transforms a prompt $p$ into a vector $F_{prompt}(p)$. The fusion model returns the final fused feature $f=\text{Norm}(\text{CrossAttn}(\text{query}=F(X),\text{key}=f_{2},\text{value}=f_{2}))$, where
[…]
$\text{query}=f_{1},\text{key}=F(X),\text{value}=F(X))))$
@@
Figure 4: Scene prior generation. We mask out objects of interest in training frames belonging to a specific scene, and average the remaining pixels across frames towards a “blank” background, which is our scene prior. A 3D detector that incorporates this scene prior (Fig. 1a) achieves remarkable improvements in roadside monocular 3D detection (Table 3).
@@
Scene prior. Hung on a roadside perception infrastructure, the camera is fixed. Hence, the scene geometry in the captured video frames can be assumed to be stable. This inspires us to derive a scene prior from these frames. To generate a scene prior, we mask out objects (by running a 2D detector on the training frames belonging to a specific scene), and then average the remaining pixels across frames to obtain a “blank” background image for this scene, which is the scene prior (Fig. 4). In this work, we simply concatenate the scene prior with the RGB frame as a six-channel input to the 3D detector. We study the affects of using different scene priors generated in daytime and nighttime, respectively; results in Table 4 show that using either remarkably enhances performance.

### v4

@@ Abstract
Figure 1: A summary of our work. (a) To approach roadside monocular 3D detection, we introduce a novel detector design Promptable 3D detector (Pro3D) that exploits 2D detections to prompt the 3D detector for 3D prediction. Pro3D can exploit any 2D detector and 3D detector with minimal modificaitons. (b) The Pro3D design is motivated by the observation that a simple 2D detector DINO [zhang2022dino] significantly outperforms state-of-the-art roadside monocular 3D detectors (e.g., BEVHeight [yang2023BEVHeight] and BEVSpread [wang2024bevspread]) w.r.t 2D metrics (on the DAIR-V2X-I dataset). This implies that 2D detection is an “easier” task than monocular 3D detection – with a trained 2D detector, training 3D detector can be “simplified” by learning to lift 2D detections to the 3D space. Moreover, as roadside camera pose is fixed, we derive a backbround image as scene prior, incorporating which remarkably boosts 3D detection performance. (c) As a summary of results, Pro3D significantly outperforms prior works on the DAIR-V2X-I benchmark. Refer to Table 1 and 2 for comprehensive results.
@@ 1 Introduction
Roadside monocular 3D detection is a practical yet challenging task that requires, over a single RGB frame, detecting objects of predefined classes (e.g., vehicle and pedestrian) and predicting their 3D properties such as the locations, shape, and orientations in bird’s-eye-view (BEV) [yang2023BEVHeight]. It has broad applications [yu2022dair, ye2022rope3d] in vehicle-vehicle communication [chen2019cooper, li2022v2x], vehicle-infrastructure cooperative perception [arnold2020cooperative], and intelligent traffic control [rauch2012car2x, wang2020v2vnet].
@@
Status quo. Roadside monocular 3D detection is an ill-posed problem as it requires inferring 3D information from 2D cues. Fortunately, camera pose is fixed on roadside infrastructure, making it possible to learn from data to infer 3D properties from a 2D image. To foster the research of roadside monocular 3D detection, the community has contributed large-scale benchmark datasets, such as DAIR-V2X-I [yu2022dair] and Rope3D [ye2022rope3d]. Existing methods train sophisticated deep neural networks in an end-to-end fashion [li2022bevformer, li2023BEVDepth, yang2023BEVHeight, wang2024bevspread]. For monocular 3D detection, a notorious difficulty is precisely estimating depth for objects [welchman2004human, rushton2009observers, gupta2023far3det]. To mitigate this difficulty, prior works propose to convert depth estimation to object height estimation and exploits camera parameters to infer object depth [yang2023BEVHeight], or consider scene geometry to carefully pool features for better 3D prediction [wang2024bevspread].
@@
Motivation. We observe that the state-of-the-art roadside 3D detectors BEVHeight [yang2023BEVHeight, yang2023bevheight++] and BEVSpead [wang2024bevspread] significantly underperform 2D detectors (e.g., DINO [zhang2022dino]) w.r.t 2D detection metrics, as shown by the precision-recall curves on the DAIR-V2X-I dataset [yu2022dair] in Fig. 1b. Here, we train the 2D detector DINO [zhang2022dino] on this dataset, using the generated 2D bounding boxes by projecting 3D cuboid annotations onto the image plane. To evaluate BEVHeight w.r.t 2D detection metrics, we project its 3D detections onto the image plane as its 2D detections. This comparison has several important implications: (1) 2D detection is an “easier” task than monocular 3D detection, and (2) the fixed camera pose or scene geometry provides an informative prior that greatly simplifies 2D detection. The results motivate us to exploit 2D detections and scene priors to facilitate 3D detector training, leading to our novel detector design (Fig. 1a), Promptable 3D Detector (Pro3D).
@@
Technical insights. Training a 3D detector involves optimizing multiple objectives [peri2022towards], such as 2D box coordinate regression, depth estimation, orientation regression, and object classification. Intuitively, optimizing them all together might cause the missing detection of hard objects as an attempt to handle the difficulty in regressing coordinates, depth and orientation. In contrast, training a 2D detector is simpler that incorporates fewer loss terms, i.e., only a 2D box coordinate regression loss and a classification loss. This observation leads to our core insight that exploits a 2D detector to facilitate 3D detector training (Fig. 1a). Intuitively, 2D detections pinpoint the target objects in the image and the 3D detector can be thought of as learning to lift them to 3D. From a contemporary view, our approach uses a 2D detector to prompt 3D detector, hence we name our framework Promptable 3D Detector (Pro3D). Moreover, as the camera is fixed in roadside perception, we derive an “empty” scene background as a scene prior from training frames (Fig. 4). We incorporate the scene prior as a part of input to the 3D detector. Extensive experiments validate the effectiveness of the scene prior, and importantly, the design of Pro3D.
@@
    To better solve roadside monocular 3D detection, we present an embarrassingly simple method to obtain a scene prior: detecting and removing objects in frames, and averaging the pixels across frames towards an “empty” scene background as the scene prior. Incorporating the scene prior greatly enhances 3D detection.
@@ 2 Related Work
Roadside monocular 3D detection can enlarg the range of perception with a camera hung on a high infrastructure [yang2023BEVHeight], serving as an important component in smart transportation and smart city by To study this problem, the community established large-scale datasets such as DAIR-V2X-I [yu2022dair] and Rope3D [ye2022rope3d]. As the roadside camera is fixed, recent methods exploit this fact to approach this problem [jinrang2024monouni, shi2023cobev, yang2023BEVHeight, yang2023bevheight++, wang2024bevspread]. For example, BEVHeight [yang2023BEVHeight] trains a model to regress towards object height, which is then converted to depth with camera extrinsics. BEVSpread [wang2024bevspread] further considers the position approximation error in the voxel pooling process based on BEVHeight. Differently, our work exploits a 2D detector (Fig. 1b) to produce 2D detections, using which to facilitate 3D detection (Fig. 1a). Further, we derive a scene prior by exploiting the fixed camera pose, and incorporate the scene prior to enhance 3D detection. Our methods can serve as a plug-in module to improve existing methods (Table 1).
@@
2D object detection aims to detect objects of predefined categories in a 2D image [felzenszwalb2009object, lin2014coco]. Prevailing approaches train deep neural networks, which typically consist of a backbone to encode input images, and a detector-head for box regression and classification [NIPS2015_fasterRCNN, redmon2016you, liu2016ssd, wang2022yolov7]. Recently, using the transformer architecture significantly improves 2D detection [detr2020carion, zhu2020deformable, zhang2022dino]. In this work, we explore how to leverage a 2D detector [zhang2022dino, wang2022yolov7] to train better 3D detectors on roadside scenes. It is important to note that any 2D detectors can be used in our framework Pro3D.
@@
Figure 2: We study three methods for encoding and fusing 2D detection prompts. Design (a) concatenates feature maps extracted by the 2D detector and the 3D detector’s backbone. Design (b) extracts a feature vector based on a 2D detection’s coordinates, encodes it through a prompt encoder, and attentively fuses it with the feature map extracted by the 3D detector’s backbone. Design (c) encodes a 2D detection, a 5-dim vector (coordinates $x$ and $y$, object width $w$ and height $h$, and the predicted class label) as the prompt, and attentively fuses the encoded detection with the feature map of the 3D detector’s backbone. Somewhat surprisingly, the third performs the best (Table 5)!
@@
Prompting, enabled by large language models (LLMs), is first recognized in Natural Language Processing (NLP) [brown2020language, sanh2021multitask, wei2022emergent]. Given an LLM, using a text prompt can instruct the LLM to produce desired text outputs. Inspired by this, the computer vision community endeavors to train large visual/multimodal models that can take text prompts to generate, restore and edit images [ramesh2021zero, liu2023explicit, nichol2021glide, Sun2024alpha]. Some works use images as prompts to accomplish vision tasks such as segmentation and inpainting [bar2022visual, wang2023images]. Beyond visual and text prompts, some recent works use more diverse information [wang2025attention] as prompts to guide models to produce desired output. For example, the Segment Anything Model (SAM) accepts bounding boxes, segmentation masks, and points as prompts towards interactive segmentation [kirillov2023segment]. In this work, we propose the promptable 3D detector design (Pro3D), which trains a 3D detector that accepts 2D detections as prompts. The instantiation of the prompt can be predicted box coordinates and class labels.
@@ 3 Promptable 3D Detector
Problem definition. Roadside monocular 3D detection requires detecting objects in a 2D RGB image and predicting their 3D information $(x,y,z,w,h,l,yaw)$, where $(x,y,z)$ is the object’s center location, $(w,h,l)$ denotes its width, height and length of the cuboid capturing the object, and yawyaw denotes its orientation on the BEV plane.
@@ 3.1 The Proposed Framework: Pro3D
Pro3D has two core modules to exploit 2D detections to facilitate 3D detection (Fig. 1a): a prompt encoder and a fusion module. The former encodes 2D detections as new features to prompt the 3D detector. The latter attentively fuses the encoded prompts with the 3D detector backbone for 3D detection. The 3D detector head makes 3D predictions based on the fused features. We elaborate on the two modules below, followed by scene prior generation.
@@
Remark 1: Using 2D detections as prompts yields better 3D detection performance than using features. To explain this, we first note different features extracted from a 2D detector: (1) global feature maps, and (2) local features at a specific position of the feature maps. Global feature maps contain all information from the input image, beneficial to object localization but lacking salient signals for semantic prediction. In contrast, local features can be informative for semantic prediction but lack spatial information for precise localization. Different from these features, the simplistic 2D detections (consisting of box coordinates, object sizes, and predicted labels) can well localize target objects in the input image and explicitly deliver their well-predicted labels. By incorporating the 2D detections, a 3D detector can be trained to lift them to the 3D space towards 3D detections. This simplifies the task of roadside monocular 3D detection and facilitates 3D detector training.
@@
Remark 2: Different from existing methods, our Pro3D explicitly leverages 2D detections. Leveraging 2D box annotations to facilitate 3D detector training is a de facto practice. A similar work to ours is MV2D [wang2023object], which trains a 3D detector from scratch for joint 2D and 3D detection. Specifically, MV2D fuses RoI features extracted from the 2D detector with those of the 3D detector’s backbone. This is similar in spirit to our explored baseline method (Fig. 2b). But our final version of Pro3D explicitly leverages 2D detections instead of their features. Extensive experiments show that Pro3D significantly outperforms MV2D (Table 1). Importantly, owing to the modular design, Pro3D can readily exploit any 2D detectors – using more advanced ones yields better 3D detection performance (Fig. 7).
Fusion module fuses the encoded prompt and the feature maps of the 3D detector backbone (Fig. 2). If the prompt is represented by a feature map extracted from the 2D detector, a straightforward fusion method is to concatenate these feature maps (Fig. 2a), as widely done in the literature [long2015fully, pinheiro2016learning]. Our final method fuses feature vectors which encode 2D detections as prompts. In this work, we implement a Transformer network [vaswani2017attention, kirillov2023segment] to attentively fuse the feature vectors with feature maps of the 3D detector backbone (Fig. 3). Concretely, for an input image $X$ which goes through the 3D detector backbone $F$, its feature map is denoted as $F(X)$. The prompt encoder $F_{prompt}$ transforms a prompt $p$ into a vector $F_{prompt}(p)$. The fusion model returns the final fused feature $f=\text{Norm}(\text{CrossAttn}(\text{query}=F(X),\text{key}=f_{2},\text{value}=f_{2}))$, where
$f_{2}=\text{Norm}($ $\text{MLP}(\text{CrossAttn}(\text{query}=f_{1},$
$\text{key}=F(X),\text{value}=F(X))))$
@@
Scene prior. Hung on a roadside perception infrastructure, the camera is fixed [zhao2021camera]. Hence, the scene geometry in the captured video frames can be assumed to be stable. This inspires us to derive a scene prior from these frames. To generate a scene prior, we mask out objects (by running a 2D detector on the training frames belonging to a specific scene), and then average the remaining pixels across frames to obtain an “empty” background image for this scene, which is the scene prior (Fig. 4). In this work, we simply concatenate the scene prior with the RGB frame as a six-channel input to the 3D detector. We study the affects of using different scene priors generated in daytime and nighttime, respectively; results in Table 4 show that using either remarkably enhances performance.
Figure 4: Scene prior generation. We mask out objects of interest in training frames belonging to a specific scene, and average the remaining pixels across frames towards an “empty” background, which is our scene prior. A 3D detector that incorporates this scene prior (Fig. 1a) achieves remarkable improvements in roadside monocular 3D detection (Table 3).
3.2 Discussions on 2D Detector Training
As Pro3D explicitly exploits a 2D detector (Fig. 1a), in principle, it can use any off-the-shelf powerful 2D detectors [Ma2023LongTailed3D]. Below we make three remarks on this matter.
Techniques of 2D detection are well established. 2D detection is more sufficiently explored than monocular 3D detection, as the former has been a fundamental problem in computer vision whereas the latter is relatively new. There are many excellent 2D detectors, such as FasterRCNN [NIPS2015_fasterRCNN], YOLO [bochkovskiy2020yolov4, wang2022yolov7], DETR [detr2020carion] and DINO [zhang2022dino]. In this work, we use the well-known transformer-based detector DINO [zhang2022dino]. We also test other 2D detectors in experiments.
More 2D detection datasets are publicly available, while the relatively new problem of monocular 3D detection has less datasets which are also small in scale. For example, the 2D detection dataset COCO [lin2014coco] (published in 2014) annotates 330K images with 2D boxes for 80 classes, whereas the 3D detection dataset nuScenes [caesar2020nuscenes] (published in 2020) “only” annotates 144K images with 3D cuboids for 23 classes. The vast amounts of 2D annotated data allows pretraining stronger 2D detectors, using which Pro3D yields more remarkable 3D detection performance gains (Fig. 7).
Joint training vs. stage-wise training. For 3D detection, most methods jointly train the 3D detector and a 2D detector by incorporating 2D annotations in an auxiliary loss [li2023BEVDepth, liu2022learning, park2021pseudo, simonelli2019disentangling, wang2022mv, wang2021fcos3d, zhang2021objects, guo2021liga, yang2023bevformer, wang2023object, luo2023latr]. Our Pro3D is a stage-wise training method, which turns out to outperform joint training (Table 6). We conjecture this is due to two reasons. First, 2D detector training is better explored with mature techniques than 3D detector training, so stage-wise training makes use of such techniques of 2D detector training. Second, joint training must strike a balance between various loss functions such as 3D cuboid regression, rotation regression and so on; yet, optimizing over all losses inevitably sacrifices 2D detection which is actually much easier than 3D detection. It is worth noting that stage-wise training allows easy leverage of external 2D datasets for pretraining a stronger 2D detector, exploiting which helps our Pro3D achieve even better performance (Fig. 7).
Figure 5: Visual comparison between the state-of-the-art method BEVHeight [yang2023BEVHeight] and our Pro3D. Results show that Pro3D can (1) make better orientation predictions than BEVHeight (column-1), and (2) detect objects (missed by BEVHeight) that are too small in size (column-2), heavily occluded (column-3), and in the far field (column-4).